# The Use of Large Language Models (LLM) for Cyber Threat Intelligence (CTI) in Cybercrime Forums

### v1

@@ 5 Conclusion
Despite these issues, OpenAI’s GPT-3.5-turbo model [7] was able to produce exceptional results across our study of 500 events. This indicates language models are currently a scalable and cost-effective solution for identifying relevant CTI data and could have exceptional potential when paired with state-of-the-art (SOTA) models for reporting data and fine-tuning results. Leveraging SOTA models, such as Claude 3.5 [2] Sonnet or GPT-4o [10, 9, 8], has the potential to further improve results or allow for categorization of even more complex data. This represents a significant opportunity for further study.
@@ Acknowledgments
This study was supported by a mini-grant from the Human-Centric Cybersecurity Partnership (HC2P).

### v3

@@ 5 Conclusion
Despite these issues, OpenAI’s GPT-3.5-turbo model [8] was able to produce exceptional results across our study. This indicates language models are currently a scalable and cost-effective solution for identifying relevant CTI data and could have exceptional potential when paired with state-of-the-art (SOTA) models for reporting data and fine-tuning results. Leveraging SOTA models, such as Claude 3.5 [2] Sonnet or GPT-4o [11, 10, 9], has the potential to further improve results or allow for categorization of even more complex data. This represents a significant opportunity for further study.
@@ Acknowledgments
This study was supported by a mini-grant from the Human-Centric Cybersecurity Partnership (HC2P). This is the second version of the manuscript. From the first version to the second one, additional data was coded to overcome heavily unbalanced variables. The manuscript was then updated with this new data, along with recall and precision metrics.